# paper1.bib

### v1

@@ 3 Subspace Newton method
$p^{N}_{k}(\hat{x}_{k})=-\beta_{k}\left(-\hat{g}_{k}-\frac{\hat{g}_{k}^{T}Hq_{k-1}}{q_{k-1}^{T}Hq_{k-1}}q_{k-1}\right),$
so we recover the method of conjugate gradients equipped with exact line search.
4 Quasi-Newton subspace approximation
To recover the $k$-th subspace Newton step from $x\in\mathcal{K}_{k}(c,H)$ to $\hat{x}_{k+1}$ as presented in the previous section, we require knowledge on how $H$ acts on $K_{k+1}(c,H)$. In particular we must have sufficient information to calculate the Newton scaling of the conjugate vectors $q_{0},\dots,q_{k-1}$ with respect to $x$ as well as the scalars $\beta_{k},\gamma_{k-1}$.
Suppose instead we simply have a quasi-Newton approximation $B_{k}\succ 0$ such that $B_{k}q_{i}=Hq_{i}$ for $i=0,\dots,k-1$. Then the only term that is unknown is $\hat{g}_{k}^{T}H\hat{g}_{k}$, which must be replaced by the approximation $\hat{g}_{k}^{T}B_{k}\hat{g}_{k}$.
If such an approximation is used, then we may construct a step $\tilde{p}_{k}(x)=p_{k-1}^{N}(x)+\delta_{k}\hat{q}_{k}$ by solving the system
$\left(\!\begin{array}[]{ccccc}\hat{g}_{k}^{T}B_{k}\hat{g}_{k}&\hat{g}_{k}^{T}Hq_{k-1}\\ q_{k-1}^{T}H\hat{g}_{k}&q_{k-1}^{T}Hq_{k-1}\end{array}\!\right)\left(\!\begin{array}[]{c}\tilde{\beta}_{k}\\ \tilde{\gamma}_{k-1}\end{array}\!\right)=\left(\!\begin{array}[]{c}-\hat{g}_{k}^{T}\hat{g}_{k}\\ 0\end{array}\!\right),$
so that
$\left(\!\begin{array}[]{c}\tilde{\beta}_{k}\\ \tilde{\gamma}_{k-1}\end{array}\!\right)=-\frac{\hat{g}_{k}^{T}\hat{g}_{k}}{\hat{g}_{k}^{T}B_{k}\hat{g}_{k}q_{k-1}^{T}Hq_{k-1}-(\hat{g}_{k}^{T}Hq_{k-1})^{2}}\left(\!\begin{array}[]{ccccc}q_{k-1}^{T}Hq_{k-1}\\ -q_{k-1}^{T}H\hat{g}_{k}\end{array}\!\right)=\delta_{k}\left(\!\begin{array}[]{c}\beta_{k}\\ \gamma_{k-1}\end{array}\!\right),$
for
$\delta_{k}=\frac{\hat{g}_{k}^{T}H\hat{g}_{k}q_{k-1}^{T}Hq_{k-1}-(\hat{g}_{k}^{T}Hq_{k-1})^{2}}{\hat{g}_{k}^{T}B_{k}\hat{g}_{k}q_{k-1}^{T}Hq_{k-1}-(\hat{g}_{k}^{T}Hq_{k-1})^{2}}.$
However, to do this we require explicit knowledge of the gradient $\hat{g}_{k}$. To avoid this, we will build an approximation $B_{k}$ which acts on the Krylov subspace $\mathcal{K}_{k}(c,H)$ in such a way that for any $x\in\mathcal{K}_{k}(c,H)$, solving the system of equations $B_{k}p_{k}(x)=-g(x)$ yields a step $p_{k}(x)=p_{k-1}^{N}(x)+\delta_{k}\hat{q}_{k}$ without requiring knowledge of $\hat{g}_{k}$
Assume that $B_{k}$ is given by
@@ 4 Quasi-Newton subspace approximation
where $Q_{k-1}=(q_{k-1}\ q_{k-2}\ \dots\ q_{0})$ for $q_{i}//\hat{q}_{i},i=0,\dots,k-1$ and $\sigma_{k}>0$. Then $B_{k}$ acts on $\mathcal{K}_{k}(c,H)$ by design.
@@
Let $x\in\mathcal{K}_{k}(c,H)$ and $B_{k}$ be given by (4.1). Then $B_{k}p_{k}=-g(x)$ has a unique solution given by
$p_{k}(x)$ $=\frac{1}{\sigma_{k}}\left(-\hat{g}_{k}+\frac{\hat{g}_{k}^{T}\!Hq_{k-1}}{q_{k-1}^{T}Hq_{k-1}}q_{k-1}\right)+\sum_{i=0}^{k-1}\left(-\frac{q_{i}^{T}\!g(x)}{q_{i}^{T}\!Hq_{i}}\right)q_{i}.$ (4.2)
Additionally there exists some unique $\sigma_{k}$ such that $p_{k}(x)=p^{N}_{k}(x)$.
Proof. Since $\hat{g}_{k}$ is the gradient at $\hat{x}_{k}$, it must belong to the null space of $Q_{k-1}$. Additionally, by construction the term
$q_{k}=-\hat{g}_{k}+\frac{\hat{g}_{k}^{T}\!Hq_{k-1}}{q_{k-1}^{T}Hq_{k-1}}q_{k-1}$
is conjugate to the vectors $q_{0},\dots,q_{k-1}$ with respect to $H$ and belongs to $K_{k+1}(c,H)$. Therefore $B_{k}q_{k}=\sigma_{k}\hat{g}_{k}$.
Additionally, given that
$p_{k-1}^{N}(x)=\sum_{i=0}^{k-1}\left(-\frac{q_{i}^{T}\!g(x)}{q_{i}^{T}\!Hq_{i}}\right)q_{i}$
we may write $g(x)=\hat{g}_{k}-Hp_{k-1}^{N}(x)$.
Applying $B_{k}$ to (4.2) therefore yields the desired result
$B_{k}p_{k}(x)$ $=-\hat{g}_{k}+\sum_{i=0}^{k-1}\left(-\frac{q_{i}^{T}\!g(x)}{q_{i}^{T}\!Hq_{i}}\right)Hq_{i}$
$=-g(x).$
For $p_{k}(x)=p_{k}^{N}(x)$ we require that each $q_{i},i=0,\dots k$ has the Newton scaling with respect to $x$, or equivalently that $q_{k}=\sigma_{k}\hat{q}_{k}$. In other words
$\sigma_{k}=-\frac{q_{k}^{T}\!Hq_{k}}{q_{k}^{T}\!g(x)}$
is the unique value for which $B_{k}$ recovers the $k$-th subspace Newton step.
The approximation $B_{k}$ given by (4.1) is independent of the choice of $x\in\mathcal{K}_{k}(c,H)$. As $B_{k}$ acts on each $q_{0},\dots,q_{k-1}$ as $H$, solving the system $B_{k}p_{k}(x)=-g(x)$ can recover the $(k-1)$-th subspace Newton step by appropriately rescaling the conjugate vectors.
Suppose that instead of using a full basis for $\mathcal{K}_{k}(c,H)$ we limit ourselves to $P_{k-1}=(p^{N}_{k-1}(x)\ q_{k-1})$. Unlike $Q_{k-1}$ which is made of a full basis of conjugate vectors, $P_{k-1}$ requires the $(k-1)$-th subspace Newton step from $x$ to $\hat{x}_{k}$, and is therefore dependent on $x$. We may then build an approximation $B_{k}(x)$ using $P_{k-1}$ by
$B_{k}(x)=\sigma_{k}(I-P_{k-1}(P_{k-1}^{T}P_{k-1})^{-1}P_{k-1}^{T})+HP_{k-1}(P_{k-1}^{T}HP_{k-1})^{-1}P_{k-1}^{T}H$ (4.3)
which will act on a two dimensional subspace of $\mathcal{K}_{k}(c,H)$ as $H$. In particular $B_{k}(x)$ holds information on how $H$ acts on the subspace Newton step from $x$ to $\hat{x}_{k}$ and the vector $q_{k}//\hat{q}_{k}$, which is sufficient to recover the step $p_{k}(x)$.
Proposition 4.2
Let $x\in\mathcal{K}_{k}(c,H)$ and $B_{k}(x)$ be given by (4.3). Then $B_{k}(x)p_{k}(x)=-g(x)$ has a unique solution given by
$p_{k}(x)$ $=\frac{1}{\sigma_{k}}\left(-\hat{g}_{k}+\frac{\hat{g}_{k}^{T}\!Hq_{k-1}}{q_{k-1}^{T}Hq_{k-1}}q_{k-1}\right)+p_{k-1}^{N}(x).$ (4.4)
@@
which coincides with the solution calculated by the approximation $B_{k}$ given by (4.1).
Proof. By the construction given in (4.3), $B_{k}(x)$ will act on $p_{k-1}^{N}(x)$ and $q_{k-1}$ as $H$. Additionally
$q_{k}=-\hat{g}_{k}+\frac{\hat{g}_{k}^{T}\!Hq_{k-1}}{q_{k-1}^{T}Hq_{k-1}}q_{k-1}$
must be conjugate to $q_{i}$ with respect to $H$ for $i=0,\dots,k-1$ and $q_{k}//\hat{q}_{k}$.
Therefore, just as $B_{k}$ given by (4.1) acts on $q_{k}$, since $q_{k}$ in the null space of $HP_{k-1}$ and $\hat{g}_{k}$ is in the null space of $P_{k-1}$, we have that $B_{k}(x)q_{k}=\sigma_{k}\hat{g}_{k}$. Applying $B_{k}(x)$ to (4.4) yields
$B_{k}p_{k}(x)$ $=-\hat{g}_{k}+Hp^{N}_{k-1}(x)$
$=-g(x).$
Given that
$p_{k-1}^{N}(x)=\sum_{i=0}^{k-1}\left(-\frac{q_{i}^{T}\!g(x)}{q_{i}^{T}\!Hq_{i}}\right)q_{i}$
we have that the equations (4.2) and (4.4) coincide.
@@ 5 An exact Quasi-Newton method without exact line search
If we are given an arbitrary point $x\in\mathcal{K}_{k}(c,H)$ then generating the $k$-th subspace Newton step or quasi-Newton step requires knowledge of the vectors $q_{i}//\hat{q}_{i}$ for $i=0,\dots k-1$. In addition it is also vital that we have information on how $H$ acts on these vectors to recover the Newton scaling. This is in general not information that is readily available.
We propose instead an iterative learning algorithm which generates subspace quasi-Newton steps using arbitrary step sizes. After a finite number of iterations, the algorithm will generate the Newton step, which will lead to termination if a unit step is used.
The step size $\alpha_{k}$ can be chosen freely. Theoretically, if we have explicit knowledge of $H$ then even a zero step will lead to generation of the Newton step as iteratively new conjugate directions are calculated and then Newton scaled with respect to the fixed point.
Without explicit knowledge of $H$ it is vital to take a nonzero step in order to evaluate a new gradient. We may then use the difference of gradients to learn how $H$ acts on the step taken. This method is described in Algorithm 1.
Algorithm 1 Quasi-Newton subspace method
$x_{0}\in\mathbb{R}^{n},B_{0}=I,p^{N}_{-1}=0$
while $k=0,1,\dots$ do
     $g_{k}\leftarrow Hx_{k}+c$
     Solve $B_{k}p_{k}=-g_{k}$
     Choose step size $\alpha_{k}\neq 0$
     $x_{k+1}\leftarrow x_{k}+\alpha_{k}p_{k}$
     $q_{k}\leftarrow p_{k}-p^{N}_{k-1}$
     $p^{N}_{k}\leftarrow(1-\alpha_{k})p^{N}_{k-1}-\left(\frac{g_{k}^{T}q_{k}}{q_{k}^{T}Hq_{k}}+\alpha_{k}\right)q_{k}$
     if $q_{k}//p^{N}_{k}$ then
         $P=\begin{bmatrix}q_{k}\end{bmatrix}$
@@
         $P=\begin{bmatrix}p^{N}_{k}&q_{k}\end{bmatrix}$
@@
     Choose some $\sigma_{k}>0$
     $B_{k+1}=\sigma_{k}(I-P(P^{T}P)^{-1}P^{T})+HP(P^{T}HP)^{-1}P^{T}H$
     $k\leftarrow k+1$
@@
For a fixed initial point $x_{0}$, consider iteration $k$ of Algorithm 1, such that $\alpha_{i}\neq 0,\sigma_{i}>0$ for $i=0,\dots,k-1$. Let $B_{k}$ be the matrix generated by the algorithm. Then $B_{k}=B(x_{k})$ as given by (4.3). Additionally, the algorithm will terminate if and only if $\alpha_{k}=1$ for some $k\geq r$, where $r$ is the dimension of the largest Krylov subspace generated by $g_{0}$ and $H$.
Proof.
To show that the Hessian approximation $B_{k}$ generated by Algorithm 1 is exactly $B_{k}(x_{k})$ given by (4.3), it is sufficient to show that $q_{k}//\hat{q}_{k}$ and that $p_{k}^{N}=p_{k}^{N}(x_{k+1})$. We proceed by induction. For $k=0$, clearly $q_{0}=-g(x_{0})$ and $p_{0}^{N}$ is Newton scaled with respect to $x_{1}$.
Assume that $q_{k-1}//\hat{q}_{k-1}$ and $p_{k-1}^{N}=p_{k-1}^{N}(x_{k})$. Then the matrix $B_{k}=B_{k}(x_{k})$ so the search direction $p_{k}$ is exactly that given by Proposition 4.2. Therefore $q_{k}=p_{k}-p_{k-1}^{N}(x_{k})//\hat{q}_{k}$.
Additionally consider the gradient $g(x_{k+1}+p_{k}^{N})$. For $q_{i},i\leq k-1$
$g(x_{k+1}+p_{k}^{N})^{T}q_{i}$ $=(g_{k}+\alpha_{k}Hp_{k-1}^{N}(x_{k})+(1-\alpha_{k})Hp_{k-1}^{N}(x_{k}))^{T}q_{i}$
$=(g_{k}+Hp_{k-1}^{N}(x_{k}))^{T}q_{i}$
$=0.$
For $q_{k}$, we may use the fact that $q_{k}$ is conjugate to $p_{k-1}^{N}(x_{k})$ with respect to $H$ to show that
$g(x_{k+1}+p_{k}^{N})^{T}q_{k}$ $=(g_{k}+\alpha_{k}Hq_{k}-(\frac{g_{k}^{T}q_{k}}{q_{k}^{T}Hq_{k}}+\alpha_{k})Hq_{k})^{T}q_{k}$
$=(g_{k}-\frac{g_{k}^{T}q_{k}}{q_{k}^{T}Hq_{k}}Hq_{k})^{T}q_{k}$
$=0.$
For $i\geq r$, the search direction $p_{i}=p_{i}^{N}$ generated by Algorithm 1 is therefore the Newton step from $x_{i+1}$ to $x^{*}$. Termination is therefore possible if and only if $\alpha_{i}=1$ for some $i>r$.
In the case that exact line search is used at each iteration of Algorithm 1, we recover BFGS with $B_{0}=\sigma_{k}I$ at iteration $k$. In such a case it is well known that $B_{k+1}$ built by the most recent conjugate vector $q_{k}$ is sufficient to guarantee finite termination.
In the case that $\alpha_{k}=1$ always, it is also sufficient to build $B_{k+1}$ using only $q_{k}$. Additionally, when unit step is always used, the algorithm will always terminate in exactly $r$ or $r+1$ iterations, as stated in the corollary below.

### v2

@@ 3 Subspace Newton method
$p^{N}_{k}(\hat{x}_{k})=\frac{1}{\hat{\sigma}_{k}}\left(-\hat{g}_{k}+\frac{\hat{g}_{k}^{T}Hq_{k-1}}{q_{k-1}^{T}Hq_{k-1}}q_{k-1}\right)=\hat{q}_{k},$
is parallel to the search direction obtained using CG, BFGS and memoryless BFGS; Newton scaled with respect to $\hat{x}_{k}$.
Given that information is known on how $H$ acts on the whole space $\mathcal{K}_{k+1}(g_{0},H)$, it is possible to calculate the subspace Newton step to $\hat{x}_{k+1}$ using the vectors $\hat{g}_{k},q_{k-1},\dots,q_{0}$. Building on this framework, we present a quasi-Newton subspace step under the assumption that we have incomplete information on the action of $H$ on the Krylov subspace $\mathcal{K}_{k+1}(g_{0},H)$.
4 Subspace quasi-Newton methods
To recover the subspace Newton step from $x\in x_{0}+\mathcal{K}_{k}(g_{0},H)$ to $\hat{x}_{k+1}$ as presented in the previous section, we require knowledge on how $H$ acts on the full subspace $K_{k+1}(g_{0},H)$. In particular sufficient information is needed to calculate the Newton scaling of the conjugate vectors $q_{0},\dots,q_{k-1}$ with respect to $x$ as well as the scalars $\beta_{k},\gamma_{k-1}$. Suppose instead we simply have a quasi-Newton approximation $B_{k}\succ 0$ such that $B_{k}q_{i}=Hq_{i}$ for $i=0,\dots,k-1$. Then the only term that is unknown is $\hat{g}_{k}^{T}H\hat{g}_{k}$, which must be replaced by the approximation $\hat{g}_{k}^{T}B_{k}\hat{g}_{k}$.
If such an approximation is used, it is possible to replace $H$ with $B_{k}$ and use the construction in Proposition 3.1 to generate a step $p_{k}(x)=q_{k}+p^{N}_{k-1}(x)$ where
$q_{k}=\frac{\hat{\sigma}_{k}}{\sigma_{k}}\hat{q}_{k}\text{for }\sigma_{k}=\frac{\hat{g}_{k}^{T}B_{k}\hat{g}_{k}q_{k-1}^{T}Hq_{k-1}-(\hat{g}_{k}^{T}Hq_{k-1})^{2}}{\hat{g}_{k}^{T}\hat{g}_{k}q_{k-1}^{T}Hq_{k-1}},$
taking into account that $B_{k}q_{i}=Hq_{i},i=0,\dots,k-1$. However, this construction requires explicit knowledge of the gradient $\hat{g}_{k}$. Additionally, $\sigma_{k}$ is undefined when $k=r$. To avoid this, we will build an approximation $B_{k}$ which acts on the Krylov subspace $\mathcal{K}_{k}(g_{0},H)$ for any $k\leq r$ in such a way that for any $x\in x_{0}+\mathcal{K}_{k}(g_{0},H)$, solving the system of equations $B_{k}p_{k}(x)=-g(x)$ yields a step $p_{k}(x)=q_{k}+p_{k-1}^{N}(x)$ where $q_{k}//\hat{q}_{k}$, without requiring knowledge of $\hat{g}_{k}$.
Lemma 4.1
For $k\leq r$, let $x\in x_{0}+\mathcal{K}_{k}(g_{0},H)$ and let $Q_{k-1}=(q_{k-1}\ q_{k-2}\ \dots\ q_{0})$ for $q_{i}//\hat{q}_{i},i=0,\dots,k-1$. Assume that $B_{k}$ is given by
@@ 4 Subspace quasi-Newton methods
where $\sigma_{k}>0$. Then $B_{k}p_{k}=-g(x)$ has a unique solution given by
$p_{k}(x)$ $=q_{k}+p^{N}_{k-1}(x)$ (4.2)
where, for $k<r$,
$q_{k}=\frac{1}{\sigma_{k}}\left(-\hat{g}_{k}+\frac{\hat{g}_{k}^{T}\!Hq_{k-1}}{q_{k-1}^{T}Hq_{k-1}}q_{k-1}\right)=\frac{\hat{\sigma}_{k}}{\sigma_{k}}\hat{q}_{k}.$
for $\hat{\sigma}_{k}$ and $\hat{q}_{k}$ given by Proposition 3.1. As a consequence, when $\sigma_{k}=\hat{\sigma}_{k}$, the solution is the subspace Newton step $p^{N}_{k}(x)$ to $\hat{x}_{k+1}$. For $k=r$, $q_{r}=\hat{q}_{r}=0$ and the solution $p_{r}(x)=p^{N}_{r-1}(x)$ is independent of the choice of $\sigma_{r}$.
Proof. Since $\hat{g}_{k}$ is the gradient at $\hat{x}_{k}$, it must be orthogonal to the space spanned by the columns of $Q_{k-1}$. Additionally, by construction the term
$q_{k}=\frac{1}{\sigma_{k}}\left(-\hat{g}_{k}+\frac{\hat{g}_{k}^{T}\!Hq_{k-1}}{q_{k-1}^{T}Hq_{k-1}}q_{k-1}\right)$
is conjugate to the vectors $q_{0},\dots,q_{k-1}$ with respect to $H$ and belongs to $K_{k+1}(g_{0},H)$. Therefore $B_{k}q_{k}=-\hat{g}_{k}$.
By Lemma 3.1, $p^{N}_{k-1}(x)$ is a linear combination of the vectors $q_{0},\dots,q_{k-1}$. Therefore $B_{k}p^{N}_{k-1}(x)=Hp^{N}_{k-1}(x)$. Since $p^{N}_{k-1}(x)$ is the subspace Newton step to $\hat{x}_{k}$, the gradient at $x$ may be rewritten as $g(x)=\hat{g}_{k}-Hp^{N}_{k-1}(x).$
Applying $B_{k}$ to (4.2) therefore yields the result
$B_{k}p_{k}(x)=-\hat{g}_{k}+Hp^{N}_{k-1}(x)=-g(x).$
As $Q_{k-1}$ has full column rank, and $\sigma_{k}>0$, $B_{k}$ is strictly positive definite by Lemma A.2 and the solution $p_{k}(x)$ is therefore unique. For the case where $\sigma_{k}=\hat{\sigma}_{k}$, the solution is exactly $p^{N}_{k}(x)=\hat{q}_{k}+p^{N}_{k-1}(x)$.
When $k=r$, the gradient $\hat{g}_{r}=0$. Therefore, by definition, $q_{r}=\hat{q}_{r}=0$. For any $\sigma_{r}>0$,
$B_{r}p^{N}_{r-1}(x)=Hp^{N}_{r-1}(x)=-g(x),$
so $p^{N}_{r-1}(x)=p_{r}(x)$.
The approximation $B_{k}$ given by (4.1) requires a full basis of $\mathcal{K}_{k}(g_{0},H)$ to calculate the step $p_{k}(x)$. In the proof of Lemma 4.1 however, it is unnecessary for $B_{k}$ to act on each basis vector $q_{0},\dots,q_{k-1}$ as $H$ individually. The key components of the proof require only that $B_{k}$ acts as $H$ on $q_{k-1}$ and $p^{N}_{k-1}(x)$.
We now show that is suffices to replace the full basis of $\mathcal{K}_{k}(g_{0},H)$ by the vectors $q_{k-1}$ and $p^{N}_{k-1}(x)$ in the construction of a Hessian approximation $B_{k}(x)$. Then $B_{k}(x)$ need only act as $H$ on a subspace of dimension at most two, regardless of the size of $\mathcal{K}_{k}(g_{0},H)$.
@@
For $k\leq r$, let $x\in x_{0}+\mathcal{K}_{k}(g_{0},H)$, let $p^{N}_{k-1}(x)$ be the subspace Newton step to $\hat{x}_{k}$ and let $q_{k-1}//\hat{q}_{k-1}$. Define
$P_{k-1}=\begin{cases}\begin{bmatrix}p^{N}_{k-1}(x)&q_{k-1}\end{bmatrix},&\text{if }\mathop{\operator@font rank}\left(\begin{bmatrix}p^{N}_{k-1}(x)&q_{k-1}\end{bmatrix}\right)=2,\\[10.0pt] \begin{bmatrix}q_{k-1}\end{bmatrix},&\text{otherwise.}\end{cases}$
Assume that $B_{k}(x)$ is given by
$B_{k}(x)=\sigma_{k}(I-P_{k-1}(P_{k-1}^{T}P_{k-1})^{-1}P_{k-1}^{T})+HP_{k-1}(P_{k-1}^{T}HP_{k-1})^{-1}P_{k-1}^{T}H.$ (4.3)
Then $B_{k}(x)p_{k}(x)=-g(x)$ has a unique solution
$p_{k}(x)$ $=q_{k}+p^{N}_{k-1}(x)$ (4.4)
@@
which coincides with the solution in Lemma 4.1 calculated by the approximation $B_{k}$ given by (4.1).
Proof. Since $k\leq r$, $q_{k-1}$ must be nonzero. Therefore $P_{k-1}$ has full column rank by definition, so $B_{k}(x)$ is well defined. Additionally, $B_{k}(x)$ is positive definite by Lemma A.2, so a solution to the equation $B_{k}(x)p_{k}(x)=-g_{k}(x)$ must exist and be unique.
Let
$q_{k}=\frac{1}{\sigma_{k}}\left(-\hat{g}_{k}+\frac{\hat{g}_{k}^{T}Hq_{k-1}}{q_{k-1}^{T}Hq_{k-1}}q_{k-1}\right).$
By construction, $q_{k}$ is conjugate to $\mathcal{K}_{k}(g_{0},H)$ with respect to $H$. By Lemma 3.1 this implies that $q_{k}^{T}Hp^{N}_{k-1}(x)=0$. Therefore $B_{k}q_{k}=-\hat{g}_{k}$. Additionally, by design $B_{k}(x)p^{N}_{k-1}(x)=Hp^{N}_{k-1}(x)$. Applying $B_{k}$ to (4.4) yields
$B_{k}(x)p_{k}(x)=-\hat{g}_{k}+Hp^{N}_{k-1}(x)=-g(x)$
which follows from $p^{N}_{k-1}(x)$ being the subspace Newton step to $\hat{x}_{k}$.
Therefore $p_{k}(x)=q_{k}+p^{N}_{k-1}(x)$ is the unique solution to the system of equations  $B_{k}(x)p_{k}(x)=-g_{k}(x)$ and coincides with the solution calculated by the approximation $B_{k}$ given by (4.1).
For any point $x$ in the affine space $x_{0}+\mathcal{K}_{k}(g_{0},H)$ we can now construct a Hessian approximation using the Newton subspace step $p^{N}_{k-1}(x)$ and a vector $q_{k-1}//\hat{q}_{k-1}$. This Hessian approximation along with the gradient yields a quasi-Newton step that is a combination of $p_{k-1}^{N}(x)$ and $q_{k}//\hat{q}_{k}$. We have shown that there is a unique value $\hat{\sigma}_{k}$ such that the resulting step is exactly the subspace Newton step $p^{N}_{k}(x)$ to $\hat{x}_{k+1}$. However, even if the resulting step is not the Newton subspace step, the point $x+\alpha p(x)$ for any step size $\alpha$ belongs to the affine space $x_{0}+\mathcal{K}_{k+1}(g_{0},H)$. We may use this property to create an iterative method which successively generates quasi-Newton steps regardless of step size.
@@ 5 An exact quasi-Newton method without exact line search
If we are given an arbitrary point $x\in x_{0}+\mathcal{K}_{k-1}(g_{0},H)$, then generating the subspace quasi-Newton step as described in Proposition 4.1 requires explicit knowledge of the direction of the subspace Newton step $p^{N}_{k-1}(x)$ from $x$ to $\hat{x}_{k}$ as well as the direction $\hat{q}_{k-1}$. In addition it is also vital that we have information on how $H$ acts on these vectors to build the Hessian approximation $B_{k}$ given by (4.3). We propose an iterative learning algorithm which generates subspace quasi-Newton steps regardless of choice of step sizes. After a finite number of iterations, the algorithm will generate the Newton step, which will lead to termination if a unit step is used. The step size $\alpha_{k}$ can be chosen freely. This method is described in Algorithm 1.
Theoretically, if we have explicit knowledge of $H$ then even a zero step will lead to generation of the Newton step as iteratively new conjugate directions are calculated and then Newton scaled with respect to the fixed point. Without explicit knowledge of $H$ it is vital to take a nonzero step in order to evaluate a new gradient. It is then possible to use a difference of gradients to learn how $H$ acts on the step taken. This is described in Lemma A.1.
Algorithm 1 An exact quasi-Newton method for solving (QP)
$x_{0}\in\mathbb{R}^{n},p^{N}_{-1}=0,\sigma_{0}>0$
$k\leftarrow 0$;
$g_{k}\leftarrow Hx_{k}+c$;
while $\|g_{k}\|\neq 0$ do
     if $k=0$ then
         $B_{k}=\sigma_{k}I$;
@@
         Choose some $\sigma_{k}>0$;
         $B_{k}=\sigma_{k}(I-P_{k-1}(P_{k-1}^{T}P_{k-1})^{-1}P_{k-1}^{T})+HP_{k-1}(P_{k-1}^{T}HP_{k-1})^{-1}P_{k-1}^{T}H$;
@@
     Solve $B_{k}p_{k}=-g_{k}$;
     Choose step size $\alpha_{k}$;
     $x_{k+1}\leftarrow x_{k}+\alpha_{k}p_{k}$;
     $g_{k+1}\leftarrow g_{k}+\alpha_{k}Hp_{k}$;
     if $\|g_{k+1}\|=0$ then
         break
     end if
     $q_{k}\leftarrow p_{k}-p^{N}_{k-1}$;
     if $\|q_{k}\|\neq 0$ then
         $p^{N}_{k}\leftarrow\left(-\frac{g_{k}^{T}q_{k}}{q_{k}^{T}Hq_{k}}-1\right)q_{k}+(1-\alpha_{k})p_{k}$;
[…]
              $P_{k}=\begin{bmatrix}p^{N}_{k}&q_{k}\end{bmatrix}$;
         else
              $P_{k}=\begin{bmatrix}q_{k}\end{bmatrix}$;
         end if
     else
         $p^{N}_{k}=(1-\alpha_{k})p^{N}_{k-1}$;
         $P_{k}=\begin{bmatrix}p^{N}_{k}\end{bmatrix}$;
     end if
     $k\leftarrow k+1$;
@@
Let $x_{0}$ be a fixed initial point and let $x_{i},i=0,\dots k$ be generated by Algorithm 1 where $\alpha_{i}$ are arbitrarily chosen for $i<k$ and $\sigma_{i}>0$ for $i\leq k$. When $k\leq r$, the search direction $p_{k}$ generated by the algorithm is the search direction $p_{k}(x_{k})$ defined in Proposition 4.1 as given by (4.4). Additionally, for $k\geq r$, the search direction $p_{k}$ generated by Algorithm 1 is precisely the Newton step so the algorithm will terminate if $\alpha_{k}=1$.
Proof. When $k<r$, to show that the search direction $p_{k+1}$ generated by Algorithm 1 is exactly $p_{k+1}(x_{k+1})$ given by (4.4), it is sufficient to show that $q_{k}//\hat{q}_{k}$ and that $p_{k}^{N}=p_{k}^{N}(x_{k+1})$. We proceed by induction. For $k=0$, clearly $\sigma_{0}q_{0}=-g(x_{0})$ and $p_{0}^{N}$ is Newton scaled with respect to $x_{1}$.
Assume that $q_{k-1}//\hat{q}_{k-1}$ and $p^{N}_{k-1}=p^{N}_{k-1}(x_{k})$, then by Proposition 4.1 $p_{k}$ is exactly $p_{k}(x_{k})$ which of the form $p_{k}=q_{k}+p^{N}_{k-1}$, therefore $q_{k}//\hat{q}_{k}$.
As the search directions generated by the algorithm coincides with those given in Proposition 4.1 for $i\leq k$, it is straightforward to show that $x_{k+1}\in x_{0}+\mathcal{K}_{k+1}(g_{0},H)$. Using the formulation for the subspace Newton step in Lemma 3.1, we have
$p_{k}^{N}(x_{k+1})$ $=-\sum_{i=0}^{k}\frac{g(x_{k+1})^{T}q_{i}}{q_{i}^{T}Hq_{i}}q_{i}$
$=-\sum_{i=0}^{k}\frac{(g_{k}+\alpha_{k}Hq_{k}+\alpha_{k}Hp_{k-1}^{N}(x_{k}))^{T}q_{i}}{q_{i}^{T}Hq_{i}}q_{i}$
$=\left(-\frac{g_{k}^{T}\!q_{k}}{q_{k}^{T}\!Hq_{k}}-\alpha_{k}\right)q_{k}-\sum_{i=0}^{k-1}\frac{g(x_{k}+\alpha_{k}p_{k-1}^{N}(x_{k}))^{T}q_{i}}{q_{i}^{T}Hq_{i}}q_{i}$
$=\left(-\frac{g_{k}^{T}\!q_{k}}{q_{k}^{T}\!Hq_{k}}-\alpha_{k}\right)q_{k}+p_{k-1}^{N}(x_{k}+\alpha_{k}p_{k-1}^{N}(x_{k}))$
$=\left(-\frac{g_{k}^{T}\!q_{k}}{q_{k}^{T}\!Hq_{k}}-\alpha_{k}\right)q_{k}+(1-\alpha_{k})p_{k-1}^{N}(x_{k})$
$=\left(-\frac{g_{k}^{T}\!q_{k}}{q_{k}^{T}\!Hq_{k}}-1\right)q_{k}+(1-\alpha_{k})p_{k},$ (5.1)
taking the conjugacy of $q_{k}$ with $p^{N}_{k-1}(x_{k}),q_{k-1},\dots,q_{0}$ into account, as well as the property of the subspace Newton step
$p^{N}_{k-1}(x_{k}+\alpha_{k}p^{N}_{k-1}(x_{k}))=\hat{x}_{k}-x_{k}-\alpha_{k}p^{N}_{k-1}(x_{k})=(1-\alpha_{k})p^{N}_{k-1}(x_{k}).$
Since $q_{k}$ is nonzero, $B_{k+1}$ is constructed by the algorithm exactly as $B_{k+1}(x_{k+1})$ is in Proposition 4.1. Therefore $p_{k+1}$ is the search direction $p_{k+1}(x_{k+1})$ given by (4.4).
As a consequence of Proposition 4.1, $p_{r}=p^{N}_{r-1}(x_{r})$ is the Newton step and $q_{r}=\hat{q}_{r}=0$. Therefore for any step $\alpha_{r}$, $p^{N}_{r}=(1-\alpha_{r})p^{N}_{r-1}$ is the Newton step from $x_{r+1}$. Proceeding again using induction, for any $k>r$, if $p^{N}_{k-1}$ is the Newton step and $q_{k}=0$, then $B_{k}$ defined by the algorithm will generate the Newton step. The algorithm will therefore terminate if $\alpha_{k}=1$.
As stated in Theorem 5.1, the search direction $p_{k}$ generated by Algorithm 1 for $k\geq n$ is precisely the Newton step. Therefore if a unit step size is chosen, the algorithm will terminate. It is however possible to terminate in exactly $r$ iterations if and only if the Hessian approximation is built using the unique value $\hat{\sigma}_{r-1}$ stated in Proposition 3.1.